A commuter bus takes 2 hours to get downtown; an express bus, averaging 25mph faster, takes 45 minutes to cover the same route. What is the average speed for the commuter bus?
Answer by htmentor(789)

A commuter bus takes 2 hours to get downtown; an express bus, averaging 25mph faster, takes 45 minutes to cover the same route. What is the average speed for the commuter bus?
============================
Let s = the speed of the commuter bus
Let d = the distance traveled by both buses
Since distance = speed x time, we can write:
d = 2s [for the commuter bus]
d = 3/4(s+25) [for the express bus; 45 min. = 3/4 hr; 25 mph faster]
Equating the RHS's gives
2s = 3/4(s+25)
Solve for s:
2s - (3/4)s = 75/4
5s/4 = 75/4
s = 75/5 = 15 mph
Ans: 15 mph